OpenStudy (anonymous):

Jose needs a rectangular section of his yard. The perimeter of the section is 27 feet, and the area is 35 square feet. Find the length and width of the section?

OpenStudy (anonymous):

Perimeter of a rectangle = 2x + 2y = 27. (1) Area = xy = 35. So x = 35/y. Sub this into (1) to find y.

OpenStudy (anonymous):

Can you write an equation for the perimeter of the section?

OpenStudy (anonymous):

the perimeter is 2(length) + 2(width) = 27.

OpenStudy (anonymous):

Using L and W, write an equation for the area of the section?

OpenStudy (anonymous):

That was the second equation, A = xy = 35.

OpenStudy (anonymous):

How do you solve for W?

OpenStudy (mathstudent55):

Let length = L Let width = W P = 2L + 2W A = LW 2L + 2W = 27 LW = 35 Solve area equation for L: L = 35/W Substitute into perimeter equation: 2(35/W) + 2W = 27 70/W + 2W = 27 Multiply both sides by W: 70 + 2W^2 = 27W 2W^2 - 27W + 70 = 0 (2W - 7)(W - 10) = 0 2W - 7 = 0 or W - 10 = 0 2W = 7 or W = 10 W = 7/2 or W = 10 Since we want the width to be smaller than the legth, we choose W = 7/2 as our answer. W = 3.5, L = 10 width = 3.5 ft; length = 10 ft